v1(3,-2,-2), v2(2,-1,-2), v3(-1,1,1) , u(0,1,h)
find all values of h such that u is a linear combination of v1, v2 and v3
1
Expert's answer
2020-03-23T14:04:30-0400

First we check if system "\{v_1, v_2, v_3 \}" is linearly independent.

"\begin{pmatrix}
 3 & -2 & -2 \\
 2 & -1 & -2 \\
 -1 & 1 & 1
\end{pmatrix} \overset{R_1 - R_2 + R_3}{=}
\begin{pmatrix}
 0 & 0 & 1 \\
 2 & -1 & -2 \\
 -1 & 1 & 1
\end{pmatrix} \overset{R_2 + 2R_3}{=}
\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 -1 & 1 & 1
\end{pmatrix} \overset{R_3 \to -R_3, R_3 + R_1 + R_2}{=}\\
=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0
\end{pmatrix}"

Thus, space of linear combinations of "v_1, v_2, v_3" includes all vectors of "R^3", so h can take on any value.